40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 17 12:55:45 2018, Koji, Summary, General, Started pumping 6x
    Reply  Tue Jul 17 18:16:38 2018, Steve, Summary, VAC, pumpdown 81 at 6 hrs 6x
       Reply  Tue Jul 17 22:25:41 2018, Koji, Summary, VAC, pumpdown 81 at 6 hrs 
          Reply  Wed Jul 18 12:49:08 2018, Steve, Summary, VAC, pumpdown 81 at 6 +4.5hrs manlGateScan.pnghanding_over_Mag.pngTGVw2auxfans_.jpg
             Reply  Wed Jul 18 17:36:50 2018, Steve, Summary, VAC, pumpdown 81 at 6 +9 hrs completed pd81completed.pngpd81@30hrs.png
                Reply  Thu Jul 19 01:56:25 2018, gautam, Summary, VAC, AUX pump shutdown 
                   Reply  Thu Jul 19 11:01:03 2018, Steve, Summary, VAC, pd81 @ 2e-5 Torr pd81@2days.png
                      Reply  Thu Jul 19 13:35:30 2018, Steve, Summary, VAC, annuloses pumped  AnsPumped.png
                         Reply  Mon Jul 23 09:58:52 2018, Steve, Summary, VAC, RGA scan at day 6 pd81-560Hz-d6.png
                            Reply  Mon Aug 6 09:34:02 2018, Steve, Update, VAC, RGA scan at day 20 pd81d20.png
                               Reply  Thu Aug 16 07:50:28 2018, Steve, Update, VAC, pumpdown 81 at day 30 pd81d30.png
                                  Reply  Thu Aug 16 14:48:14 2018, Steve, Update, VAC, why do we need a root pump? 6x
                                     Reply  Fri Nov 16 10:26:12 2018, Steve, Update, VAC, single viton O-rings Single-O-ring.png
                                  Reply  Fri Aug 24 10:51:23 2018, Steve, Update, VAC, pumpdown 81 at day 38 d38.png
                                     Reply  Thu Oct 4 08:25:50 2018, Steve, Update, VAC, rga scan pd81 at day 78 pd81d78.png
                                        Reply  Fri Oct 12 08:27:05 2018, Steve, Update, VAC, drypump forepump.png
                                           Reply  Fri Oct 12 17:37:03 2018, Steve, Update, VAC, pressure gauge choices  
                                           Reply  Mon Oct 15 13:59:42 2018, Steve, Update, VAC, drypump replaced drypump_swap.png
                                        Reply  Tue Nov 13 16:15:01 2018, Steve, Update, VAC, rga scan pd81 at day 119 pd81-d119.pngpd81-560Hz-d119.png
Message ID: 14077     Entry time: Tue Jul 17 12:55:45 2018     Reply to this: 14079
Author: Koji 
Type: Summary 
Category: General 
Subject: Started pumping 

[Steve, Koji, Gautam]

We started pumping down at ~12:15PM.

Vent finalization ~ YEND

  • The table leveling was way off. This was adjusted by the balancing weight. (Attachment 1~3)
  • The alignment of ETMY was not too much off. Just aligned it with the oplev spot on MEDM and this already made the green flashing.
  • The Green TEM00 was maximized with ITMY and ETMY. This made the PSL IR flashing.
  • The heater wires were checked. I found that one of the heater wires was touching the optical table via the cable shield. This is because the upper pins were shifted to the left side (Attachment 4&5). The pins were shifted and now all 4 cables are isolated form the table. I also checked the mutual resistance between the 4 terminals. They were measured to be isolated except two pairs that showed 4.4 Ohms and 4.0 Ohms (Attachment 6)
  • The tools were removed from the chamber. The Y arm was still flashing.
  • We closed the ETMY door.

Vent finalization ~ Vertex

  • Found the ITMX stuck. Gautam came in and showed us his black magic to release the optic...
  • This allowed us to align X arm. The green flash was found and the TEM00 flash was seen. This allowed us to see the PSL IR flash at the X end.
  • PRM Refl was aligned. SRM was aligned with the oplev.
  • The beam on the AS port was checked. The AS beam came out from the window.
  • Closed the OMC chamber.

Pumping

  • Started pumping with RP1 and RP3. (~12:15PM)

Attachment 1: IMG_5408.JPG  729 kB  Uploaded Tue Jul 17 14:23:29 2018  | Hide | Hide all | Show all
IMG_5408.JPG
Attachment 2: IMG_5400.JPG  521 kB  Uploaded Tue Jul 17 14:23:36 2018  | Show | Hide all | Show all
Attachment 3: IMG_5401.JPG  559 kB  Uploaded Tue Jul 17 14:23:41 2018  | Show | Hide all | Show all
Attachment 4: IMG_5402.JPG  1.757 MB  Uploaded Tue Jul 17 14:30:44 2018  | Show | Hide all | Show all
Attachment 5: IMG_5403.JPG  1.080 MB  Uploaded Tue Jul 17 14:30:51 2018  | Show | Hide all | Show all
Attachment 6: IMG_5404.JPG  805 kB  Uploaded Tue Jul 17 14:33:37 2018  | Hide | Hide all | Show all
IMG_5404.JPG
ELOG V3.1.3-